OpenStudy (anonymous):

Solve the exponential equation 2^x=4^x+1

OpenStudy (anonymous):

Or 2(x)=4(x+1) I

OpenStudy (anonymous):

2^x=4^(x+1) 2^X=2^2(X+1) x=2(x+1) x=2x+2 x=-2

OpenStudy (solomonzelman):

I don't know how did @mohammad.arqum get the second step?

OpenStudy (anonymous):

I USED................... . . . . . . THE POWER OF MATH...

OpenStudy (solomonzelman):

but it is \[4^{(x+1)}=2^2(x+1)\]\[4^{(x+1)}=4 \times (x+1)\] true or false?

OpenStudy (solomonzelman):

This is supposed to be\[2(x)=4(x+1)\]\[2x=4x+4\]\[2x\color{blue} {-4x } =4x+4\color{blue} {-4x } \]\[-2x=4\] \[......\]

OpenStudy (anonymous):

i cant understand what u are writing its all scrambled and stuff \[2(x)=4(x+1)\]\[2x=4x+4\]\[2x\color{blue} {-4x } =4x+4\color{blue} {-4x } \]\[-2x=4\] \[......\]

OpenStudy (solomonzelman):

The question is without any exponents, the simplest algebra. expand the parenthesis on both sides, subtract 4x from both sides, and then the final step that the user should know how to do. what you wrote was amorphous...

OpenStudy (solomonzelman):

Go over your work, you should be able to see your mistake, if not, then I would suggest use this site as an asker, or someone that needs help, rather then helper.

OpenStudy (anonymous):

NNOOOO......the answer i wrote was correct u're mixing up the question the question was

OpenStudy (solomonzelman):

you didn't write the question correctly, sir, I don't see that the person who asked the question, mentioned any exponent in his problem? you are doing something very complicated, instead of taking the BASIC APPROACH. Your answer is correct though, apologize for this, but why are you doing that way?

OpenStudy (solomonzelman):

He wrote it 1st with exponents, but the re-wrote it.

OpenStudy (anonymous):

2^x=4^(x+1) this was the question as the questioner has told there are exponents the rewrite was supposed to define the question more to remove any misunderstandings telling that (x+1) was a power not that x was only a power with 1 added later

OpenStudy (solomonzelman):

If you understand "2(x)=4(x+1)" as a problem with exponents, then discussing anything any farther, is a waste of time.

OpenStudy (anonymous):

FACEPALM leave it wud ya

OpenStudy (solomonzelman):

okay nvm, I am not wasting my time on this any more.